In eq. E=1/(4πε_o )∫ρ_v R" " /R^3 dv' , let ρ_v to be known. Explain in detail how you would numerically evaluate the vector integral to obtain E.


1
Expert's answer
2020-08-10T19:57:52-0400

As per the question,

"E=\frac{1}{(4πε_o )}\int \frac{\rho_v R}{R^3}dv"


"=\frac{1}{4\pi \epsilon_o}\int \frac{\rho_v R 4\pi dv}{4\pi R^3}"


"=\frac{1}{4\pi \epsilon_o}\int \frac{\rho_v R 4\pi dv}{v}"


"=\frac{\rho_o R\times 4 \pi}{4\pi \epsilon_o} \ln(v)+c"


"=\frac{\rho_o R}{\epsilon_o}\ln (v)+c"

Hence, the required electric field will be calculated from the above.